OpenStudy (anonymous):

Medal will be given!! When 5 is added to 2 times a number, the result is 45. Find the number. A. 20 B. 45 C. 5 D. 2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

x = 20

OpenStudy (anonymous):

5 + 2x =45 2x = 40 x = 20

OpenStudy (anonymous):

Thanks! Here's another one. Select the algebraic equation that correctly represents the given sentence. Jane currently has $3,360 in her savings account. Prior to getting paid, she had $1,230 in her savings account. How much was Jane's paycheck? A. $1,230 + $3,360 = x x = $4,590 B. $1,230 + x = $3,360 x = $2,130 C. $1,230x = $3,360 x = $2.73 D. none of the above

OpenStudy (triciaal):

amount she got paid = x 3360 = x +1230 solve for x

OpenStudy (anonymous):

How do I solve for x? @triciaal

OpenStudy (triciaal):

find what x =

OpenStudy (triciaal):